To make 2.000 gal of a powdered sports drink, a group of students measure out 2.000 gal of water with 500.-mL, 50.-mL, and 5-mL graduated cylinders. Show how they could get closest to 2.000 gal of water, using these cylinders the fewest times.
1
Expert's answer
2021-09-11T00:04:52-0400

1 gal =3785.412 ml




2 × 3785.412 = 7570.824ml




15 × 500 = 7500




7570.824-7500 = 70.824




1 × 50 = 50




70.824-50 = 20.824




5 × 4 = 20




20.824 -20 = 0.824ml




= 15 (500ml), 1(50ml), 4(5ml)